# Elliptic Quantum Toroidal Algebra $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor}),$ Vertex Operators and $L$-operators Hitoshi Konno and Andrey Smirnov

### v1

@@ 3.1 The instanton moduli space ${\cal M}(n,r)$
${\cal M}(n,1)\cong\mathrm{Hilb}^{n}({\mathbb{C}}^{2}).$
@@
${\cal M}(n,r)^{A}=\coprod_{n_{1}+\cdots+n_{r}=n}\mathrm{Hilb}^{n_{1}}({\mathbb{C}}^{2})\times\cdots\times\mathrm{Hilb}^{n_{r}}({\mathbb{C}}^{2}).$
Let us consider the case $r=1$, the Hilbert scheme ${\cal H}_{n}=\mathrm{Hilb}^{n}({\mathbb{C}}^{2})$, $A={\mathbb{C}}^{\times}$. We denote the coordinate on $A$ by $u$ such that
@@
The fixed points ${\cal H}_{n}^{T}={\cal H}_{n}^{A}$ is a finite set labeled by partitions of $n$. Let
${\cal P}_{n}=\{\ \lambda\in{\cal P}\ |\ |\lambda|=n\ \}.$ (3.1)
We regard $\lambda\in{\cal P}_{n}$ as a Young diagram with $n$ boxes. For a box $\square=(i,j)\in\lambda$, we deine
@@
$\rho_{\square}=c_{\square}-\epsilon h_{\square}$
@@
Let us consider the following presentation of the equivariant $\mathrm{K}$-theory of ${\cal H}_{n}$.
$\mathrm{K}_{T}({\cal H}_{n})={\mathbb{Z}}[x_{1}^{\pm},\cdots,x_{n}^{\pm},t_{1}^{\pm},t_{2}^{\pm}]^{\mathfrak{S}_{n}}/R,$
where $\mathfrak{S}_{n}$ denotes the symmetrization in $x_{a}$’s, and $R$ the ideal of Laurent polynomials vanishing at all fixed points in ${\cal H}_{n}^{T}$. We often loosely use $x_{a}$ as $x_{\iota(a)}$ for $a\in\lambda$ and vice versa. For $\square=(i,j)\in\lambda$, we set
@@
The restriction of a $\mathrm{K}$-theory class $f(x_{1},\cdots,x_{n},t_{1},t_{2})$ to a fixed point labeled by $\lambda\in{\cal P}_{n}$ is given by
@@
Here $i_{\lambda}\ :\ \lambda\ \to\ {\cal H}_{n}$ denotes the canonical inclusion of a fixed point.
Let ${\cal V}$ be the rank $n$ tautological bundle on ${\cal H}_{n}$. We present ${\cal V}$ as
${\cal V}=x_{1}+\cdots+x_{n}$
regarding $x_{1},\cdots,x_{n}$ as the Chern roots of ${\cal V}$. The line bundle
${\cal O}(1)=\det{\cal V}=x_{1}\cdots x_{n}$
generates the Picard group $\mbox{Pic}({\cal H}_{n})={\mathbb{Z}}$.
In general, fixed points in ${\cal M}(n,r)$ are lebeled by a $r$-tuple partition ${\boldsymbol{\lambda}}=(\lambda^{(1)},\cdots,\lambda^{(r)})$ with $|{\boldsymbol{\lambda}}|=\sum_{i=1}^{r}|\lambda^{(i)}|=n$. We denote the Chern roots of the rank $n$ tautological bundle ${\cal V}$ on ${\cal M}(n,r)$ by ${\bf x}=(x_{1},\cdots,x_{n})$ and a coordinate of $A=({\mathbb{C}}^{\times})^{r}$ by ${\bf u}=(u_{1},\cdots,u_{r})$. We define a canonical ordering on the $n$ boxes of ${\boldsymbol{\lambda}}$ by extending the one for each partition $\lambda^{(i)}$ with adding the following condition, for $i<j$
@@
Note that this additional condition is realized as a condition $\rho_{a}<\rho_{b}$ if one extends the definition of $\rho_{a}$ to
$\rho_{a}:=c_{a}-\epsilon h_{a}+u_{i}\qquad\mbox{for}\ a\in\lambda^{(i)}$ (3.2)
and assume $|u_{1}|\ll\cdots\ll|u_{r}|$. In the following we use this convention for ${\boldsymbol{\lambda}}$.
3.2 Elliptic stable envelopes
The elliptic stable envelopes for $\mathrm{E}_{T}({\cal H}_{n})$ was constructed in [44]. Let $T^{1/2}\in\mathrm{K}_{T}({\cal H}_{n})$ be a polarization of ${\cal H}_{n}$ satisfying
$T{\cal H}_{n}=T^{1/2}+\hbar(T^{1/2})^{\vee}\in\mathrm{K}_{T}({\cal H}_{n}).$
The term $\hbar(T^{1/2})^{\vee}$ is called the opposite polarization. For $\lambda\in{\cal P}_{n}$, let us set
$S^{Ell}_{\lambda}(x_{1},\cdots,x_{n},u){}$
$:=\frac{\prod_{a,b\in\lambda\atop\rho_{a}+1<\rho_{b}}\theta(t_{1}x_{a}/x_{b})\prod_{a,b\in\lambda\atop\rho_{a}+1>\rho_{b}}\theta(t_{2}x_{b}/x_{a})\prod_{a\in\lambda\atop\rho_{a}\leq 0}\theta(x_{a}/u)\prod_{a\in\lambda\atop\rho_{a}>0}\theta(\hbar u/x_{a})}{\prod_{a,b\in\lambda\atop\rho_{a}<\rho_{b}}\theta(x_{a}/x_{b})\theta(\hbar x_{a}/x_{b})}.$ (3.3)
@@ 3.2 Elliptic stable envelopes
$W^{Ell}_{T^{1/2}}({\bf t};x_{1},\cdots,x_{n},u,\mathfrak{z})=(-1)^{\kappa_{{\bf t}}}\phi\Big{(}\frac{x_{r}}{u},\mathfrak{z}^{n}(t_{1}t_{2})^{\mathsf{v}_{r}}\Big{)}\prod\limits_{e\in{\bf t}}\,\phi\Big{(}\dfrac{x_{h(e)}\varphi^{\lambda}_{t(e)}}{x_{t(e)}\varphi^{\lambda}_{h(e)}},\mathfrak{z}^{\mathsf{w}_{e}}(t_{1}t_{2})^{\mathsf{v}_{e}}\Big{)}$
@@
where the product runs over the edges of the tree ${\bf t}$ and $h(e)\in\lambda$, $t(e)\in\lambda$ denote the head and tail box of the edge $e$. And, $\kappa_{{\bf t}},\mathsf{w}_{e},\mathsf{v}_{e}\in\mathbb{Z}$ are certain integers computed from the tree in a combinatorial way, we refer to Sections 4.2-4.5 in [44] for definitions of these integers. We also identify the dynamical parameter $\mathfrak{z}$ with the Kähler parameter in $\mathrm{E}_{T}({\cal H}_{n})$[1] in the same way as in [29]. Finally, let $\Upsilon_{\lambda}$ be the set of $\lambda$-trees without $\mathsf{L}$-shaped subgraphs, see section 4.6 in [44]. Then, we have:
@@
[44] The elliptic stable envelope of a fixed point $\lambda\in{\cal H}_{n}^{A}$ is given by
$\mathrm{Stab}_{\mathfrak{C},T^{1/2}}(\lambda;\mathfrak{z})=\mathrm{Sym}\left(S^{Ell}_{\lambda}(x_{1},\cdots,x_{n},u)\sum_{\delta\in\Upsilon_{\lambda}}W^{Ell}_{T^{1/2}}({\bf t}_{\delta};x_{1},\cdots,x_{n},u;\mathfrak{z})\right),$ (3.4)
@@
The elliptic stable envelope for $\mathrm{E}_{T}({\cal M}(n,r))$ is constructed by taking the shuffle product[2] of those for the Hilbert schemes. Let $\lambda^{\prime},\lambda^{\prime\prime}$ be two partitions with $|\lambda^{\prime}|=n^{\prime},|\lambda^{\prime\prime}|=n^{\prime\prime}$ and consider the elliptic stable envelopes $\mathrm{Stab}_{T^{1/2},\mathfrak{C}^{\prime}}(\lambda^{\prime};\mathfrak{z})$ and $\mathrm{Stab}_{T^{1/2},\mathfrak{C}^{\prime\prime}}(\lambda^{\prime\prime};\mathfrak{z})$ for $\mathrm{E}_{T}({\cal H}_{n^{\prime}})$ and $\mathrm{E}_{T}({\cal H}_{n^{\prime\prime}})$ with the equivariant parameters $u_{1},u_{2}$, respectively. Here one takes $\mathfrak{C}^{\prime}=\mathfrak{C}^{\prime\prime}$ as $t_{1}/t_{2}>0$. We take the canonical ordering on the $n^{\prime}+n^{\prime\prime}$ boxes in the double partition $(\lambda^{\prime},\lambda^{\prime\prime})$ as defined in Sec.3.1. Let us consider ${\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}((\lambda^{\prime},\lambda^{\prime\prime});\mathfrak{z})$ given by
${\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}((\lambda^{\prime},\lambda^{\prime\prime});\mathfrak{z}){}$
@@
$:=\mathrm{Sym}_{\{x_{a}\}_{a\in(\lambda^{\prime},\lambda^{\prime\prime})}}\left(\begin{matrix}&\cr&\cr\end{matrix}\prod_{a\in\lambda^{\prime},b\in\lambda^{\prime\prime}}\frac{\theta(t_{1}x^{\prime}_{a}/x^{\prime\prime}_{b})\theta(t_{2}x^{\prime}_{a}/x^{\prime\prime}_{b})}{\theta(x^{\prime}_{a}/x^{\prime\prime}_{b})\theta(\hbar x^{\prime}_{a}/x^{\prime\prime}_{b})}\prod_{b\in\lambda^{\prime\prime}}{\theta(\hbar u_{1}/x_{b}^{\prime\prime})}\prod_{a\in\lambda^{\prime}}\theta(x^{\prime}_{a}/u_{2})\right.{}$
@@
×$\left.\hskip 113.81102pt\qquad\qquad\times\begin{matrix}&\cr&\cr\end{matrix}{\mathrm{Stab}}_{\mathfrak{C}^{\prime},T^{1/2}}(\lambda^{\prime};\mathfrak{z}\hbar^{-1}){\mathrm{Stab}}_{\mathfrak{C}^{\prime\prime},T^{1/2}}(\lambda^{\prime\prime};\mathfrak{z})\right).$ (3.5)
@@
Here $x_{a}^{\prime}\ (a\in\lambda^{\prime})$ and $x_{b}^{\prime\prime}\ (b\in\lambda^{\prime\prime})$ are the Chern roots for the tautological bundle on ${\cal H}_{n^{\prime}}$ and ${\cal H}_{n^{\prime\prime}}$, respectively, and we set $\{x_{a}\}_{a\in(\lambda^{\prime},\lambda^{\prime\prime})}=\{x_{a}^{\prime}\}_{a\in\lambda^{\prime}}\cup\{x_{b}^{\prime\prime}\}_{b\in\lambda^{\prime\prime}}$. Then one finds that ${\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}((\lambda^{\prime},\lambda^{\prime\prime});\mathfrak{z})$ is the elliptic stable envelope for $\mathrm{E}_{T}({\cal M}(n^{\prime}+n^{\prime\prime},2))$ with the chamber $\mathfrak{C}$ given by $|u_{1}|\ll|u_{2}|$. The formula (3.5) is called the shuffle product [2]. By abuse of notation we use $T^{1/2}$ in ${\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}((\lambda^{\prime},\lambda^{\prime\prime});\mathfrak{z})$ as $T^{1/2}{\cal M}$ i.e. the polarization satisfying
$T{\cal M}=T^{1/2}{\cal M}+\hbar(T^{1/2}{\cal M})^{\vee}\in\mathrm{K}_{T}({\cal M})$
@@
for ${\cal M}={\cal M}(n^{\prime}+n^{\prime\prime},2)$. The shuffle product keeps such assignment of polarization consistent through a construction of the elliptic stable envelopes. See the remark below. In the following, we use the same convention.
@@
By repeating the shuffle product, one obtains the elliptic stable envelopes for $\mathrm{E}_{T}({\cal M}(n,r))$. Let $\lambda^{(i)}\ (i=1,\cdots,r)$ be partitions with $|\lambda^{(i)}|=n_{i}$ and consider a $r$-tuple partition ${\boldsymbol{\lambda}}=(\lambda^{(1)},\cdots,\lambda^{(r)})$. We set $|{\boldsymbol{\lambda}}|=\sum_{i=1}^{r}n_{i}=n$.
@@
The following gives the elliptic stable envelopes for $\mathrm{E}_{T}({\cal M}(n,r))$ with the chamber $\mathfrak{C}$ being $|u_{1}|\ll\cdots\ll|u_{r}|$.
@@
$S^{Ell}_{\boldsymbol{\lambda}}({\bf x},{\bf u}):=\frac{\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}+1<\rho_{b}}\theta(t_{1}x_{a}/x_{b})\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}+1>\rho_{b}}\theta(t_{2}x_{b}/x_{a})}{\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}<\rho_{b}}\theta(x_{a}/x_{b})\theta(\hbar x_{a}/x_{b})}\prod_{i=1}^{r}\left[\prod_{a\in{\boldsymbol{\lambda}}\atop\rho_{a}>\rho_{r_{i}}}{\theta(\hbar u_{i}/x_{a})}\prod_{a\in{\boldsymbol{\lambda}}\atop\rho_{a}\leq\rho_{r_{i}}}\theta(x_{a}/u_{i})\right],$
$\sum_{{\bf t}}W^{Ell}_{T^{1/2}}({\bf t};{\bf x},{\bf u};\mathfrak{z})$
$:=\sum_{\delta^{(1)}\in\Upsilon_{\lambda^{(1)}},\cdots,\delta^{(r)}\in\Upsilon_{\lambda^{(r)}}}W^{Ell}_{T^{1/2}}({\bf t}^{(1)}_{\delta^{(1)}};{\bf x}^{(1)},u_{1};\mathfrak{z}\hbar^{-r+1})W^{Ell}_{T^{1/2}}({\bf t}^{(2)}_{\delta^{(2)}};{\bf x}^{(2)},u_{2};\mathfrak{z}\hbar^{-r+2})\cdots{}$
×$\hskip 142.26378pt\cdots\times W^{Ell}_{T^{1/2}}({\bf t}^{(r)}_{\delta^{(r)}};{\bf x}^{(r)},u_{r};\mathfrak{z})$
with $r_{i}$ being the root box of $\lambda^{(i)}$, ${\bf x}^{(i)}=(x^{(i)}_{1},\cdots,x^{(i)}_{n_{i}})$ being the Chern roots associated to the Hilbert scheme ${\cal H}_{n_{i}}$ and ${\bf t}:=({\bf t}^{(1)}_{\delta^{(1)}},\cdots,{\bf t}^{(r)}_{\delta^{(r)}})$.
Remark.  A more general formula for the elliptic stable envelope of affine $A$-type quivers is obtained in [3]. For ${\cal M}(n,r)$, this result coincides with Proposition 3.2. We note that the dynamical shifts of $\mathfrak{z}$ is controlled in [3] via the positive part of polarization bundle $\det T^{1/2}_{\rho>0,\lambda}$, see formula (13) and Definition 4.2.1 in [3]. For $\rho$ defined by (3.2) and the chamber $|u_{1}|\ll\cdots\ll|u_{r}|$, these shifts are given by subtitutions $\mathfrak{z}\to\mathfrak{z}\hbar^{-r+i}$ in the functions $W^{Ell}_{T^{1/2}}$, exactly as in the formulas of Proposition 3.2.

### v2

@@ 3.1. The instanton moduli space ${\mathcal{M}}(n,r)$
${\mathcal{M}}(n,1)\cong\mathrm{Hilb}^{n}({\mathbb{C}}^{2}).$
@@
${\mathcal{M}}(n,r)^{A}=\coprod_{n_{1}+\cdots+n_{r}=n}\mathrm{Hilb}^{n_{1}}({\mathbb{C}}^{2})\times\cdots\times\mathrm{Hilb}^{n_{r}}({\mathbb{C}}^{2}).$
Let us consider the case $r=1$, the Hilbert scheme ${\mathcal{H}}_{n}=\mathrm{Hilb}^{n}({\mathbb{C}}^{2})$, $A={\mathbb{C}}^{\times}$. We denote the coordinate on $A$ by $u$ such that
@@
The $A$-fixed points of ${\mathcal{H}}_{n}$ are labeled by partitions of $n$. Let
${\mathcal{P}}_{n}=\{\ \lambda\in{\mathcal{P}}\ |\ |\lambda|=n\ \}.$ (3.1)
We regard $\lambda\in{\mathcal{P}}_{n}$ as a Young diagram with $n$ boxes. For a box $\square=(i,j)\in\lambda$, we define
@@
$\rho_{\square}:=c_{\square}-\epsilon h_{\square}$
@@
Let us consider the following presentation of the equivariant $\mathrm{K}$-theory of ${\mathcal{H}}_{n}$.
$\mathrm{K}_{T}({\mathcal{H}}_{n})={\mathbb{Z}}[x_{1}^{\pm},\cdots,x_{n}^{\pm},t_{1}^{\pm},t_{2}^{\pm}]^{\mathfrak{S}_{n}}/R,$
where $\mathfrak{S}_{n}$ denotes the symmetrization in $x_{a}$’s, and $R$ the ideal of Laurent polynomials vanishing at all fixed points in ${\mathcal{H}}_{n}^{T}$. We often loosely use $x_{a}$ as $x_{\iota(a)}$ for $a\in\lambda$ and vice versa. For $\square=(i,j)\in\lambda$, we set
@@
The restriction of a $\mathrm{K}$-theory class $f(x_{1},\cdots,x_{n},t_{1},t_{2})$ to a fixed point labeled by $\lambda\in{\mathcal{P}}_{n}$ is given by
@@
Here $i_{\lambda}\ :\ \lambda\ \to\ {\mathcal{H}}_{n}$ denotes the canonical inclusion of a fixed point.
Let $\mathcal{V}$} be the rank $n$ tautological bundle on ${\mathcal{H}}_{n}$. We present $\mathcal{V}$} as
${\mathcal{V}}=x_{1}+\cdots+x_{n}$
regarding $x_{1},\cdots,x_{n}$ as the Chern roots of $\mathcal{V}$}. The line bundle
${\mathcal{O}}(1)=\det{\mathcal{V}}=x_{1}\cdots x_{n}$
generates the Picard group $\mbox{Pic}({\mathcal{H}}_{n})={\mathbb{Z}}$.
In general, fixed points in ${\mathcal{M}}(n,r)$ are labeled by a $r$-tuple partition ${\boldsymbol{\lambda}}=(\lambda^{(1)},\cdots,\lambda^{(r)})$ with $|{\boldsymbol{\lambda}}|=\sum_{i=1}^{r}|\lambda^{(i)}|=n$. We denote the Chern roots of the rank $n$ tautological bundle $\mathcal{V}$} on ${\mathcal{M}}(n,r)$ by ${\bf x}=(x_{1},\cdots,x_{n})$ and a coordinate of $A=({\mathbb{C}}^{\times})^{r}$ by ${\bf u}=(u_{1},\cdots,u_{r})$. We define a canonical ordering on the $n$ boxes of $\boldsymbol{\lambda}$} by extending the one for each partition $\lambda^{(i)}$ with adding the following condition, for $i<j$
@@
3.2. Elliptic stable envelopes
The elliptic stable envelopes for $\mathrm{E}_{T}({\mathcal{H}}_{n})$ was constructed in [44]. Let $T^{1/2}\in\mathrm{K}_{T}({\mathcal{H}}_{n})$ be a polarization of ${\mathcal{H}}_{n}$ satisfying
$T{\mathcal{H}}_{n}=T^{1/2}+\hbar(T^{1/2})^{\vee}\in\mathrm{K}_{T}({\mathcal{H}}_{n}).$
The term $\hbar(T^{1/2})^{\vee}$ is called the opposite polarization. For $\lambda\in{\mathcal{P}}_{n}$, let us set
$S^{Ell}_{\lambda}(x_{1},\cdots,x_{n},u)$ (3.2)
$:=\frac{\prod_{a,b\in\lambda\atop\rho_{a}+1<\rho_{b}}\theta(t_{1}x_{a}/x_{b})\prod_{a,b\in\lambda\atop\rho_{a}+1>\rho_{b}}\theta(t_{2}x_{b}/x_{a})}{\prod_{a,b\in\lambda\atop\rho_{a}<\rho_{b}}\theta(x_{a}/x_{b})\theta(\hbar x_{a}/x_{b})}{}$
$\qquad\times\prod_{a\in\lambda\atop\rho_{a}\leq 0}\theta(x_{a}/u)\prod_{a\in\lambda\atop\rho_{a}>0}\theta(\hbar u/x_{a}).{}$
@@ 3.2. Elliptic stable envelopes
$W^{Ell}_{T^{1/2}}({\bf t};x_{1},\cdots,x_{n},u,\mathfrak{z})$
$=(-1)^{\kappa_{{\bf t}}}\phi\Big{(}\frac{x_{r}}{u},\mathfrak{z}^{n}(t_{1}t_{2})^{\mathsf{v}_{r}}\Big{)}\prod\limits_{e\in{\bf t}}\,\phi\Big{(}\dfrac{x_{h(e)}\varphi^{\lambda}_{t(e)}}{x_{t(e)}\varphi^{\lambda}_{h(e)}},\mathfrak{z}^{\mathsf{w}_{e}}(t_{1}t_{2})^{\mathsf{v}_{e}}\Big{)}$
@@
where the product runs over the edges of the tree $\bf t$} and $h(e)\in\lambda$, $t(e)\in\lambda$ denote the head and tail box of the edge $e$. And, $\kappa_{{\bf t}},\mathsf{w}_{e},\mathsf{v}_{e}\in\mathbb{Z}$ are certain integers computed from the tree in a combinatorial way, we refer to Sections 4.2-4.5 in [44] for definitions of these integers. We identify the dynamical parameter $\mathfrak{z}$ with the Kähler parameter in $\mathrm{E}_{T}({\mathcal{H}}_{n})$[1] in the same way as in [29]. Finally, let $\Upsilon_{\lambda}$ be the set of $\lambda$-trees without $\mathsf{L}$ -shaped subgraphs, see section 4.6 in [44]. Then, we have:
@@
[44] The elliptic stable envelope of a fixed point $\lambda\in{\mathcal{H}}_{n}^{A}$ is given by
$\hskip 8.53581pt\mathrm{Stab}_{\mathfrak{C},T^{1/2}}(\lambda;z)$ (3.3)
$\ =\mathrm{Sym}\left(S^{Ell}_{\lambda}(x_{1},\cdots,x_{n},u)\sum_{\delta\in\Upsilon_{\lambda}}W^{Ell}_{T^{1/2}}({\bf t}_{\delta};x_{1},\cdots,x_{n},u;\mathfrak{z})\right),{}$
@@
The elliptic stable envelope for $\mathrm{E}_{T}({\mathcal{M}}(n,r))$ is constructed by taking the shuffle product[2] of those for the Hilbert schemes. Let $\lambda^{\prime},\lambda^{\prime\prime}$ be two partitions with $|\lambda^{\prime}|=n^{\prime},|\lambda^{\prime\prime}|=n^{\prime\prime}$ and consider the elliptic stable envelopes $\mathrm{Stab}_{T^{1/2},\mathfrak{C}^{\prime}}(\lambda^{\prime};\mathfrak{z})$ and $\mathrm{Stab}_{T^{1/2},\mathfrak{C}^{\prime\prime}}(\lambda^{\prime\prime};\mathfrak{z})$ for $\mathrm{E}_{T}({\mathcal{H}}_{n^{\prime}})$ and $\mathrm{E}_{T}({\mathcal{H}}_{n^{\prime\prime}})$ with the equivariant parameters $u_{1},u_{2}$, respectively. Here one takes $\mathfrak{C}^{\prime}=\mathfrak{C}^{\prime\prime}$ as $t_{1}/t_{2}>0$. We take the canonical ordering on the $n^{\prime}+n^{\prime\prime}$ boxes in the double partition $(\lambda^{\prime},\lambda^{\prime\prime})$ as defined in Sec.3.1. Let us consider ${\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}((\lambda^{\prime},\lambda^{\prime\prime});\mathfrak{z})$ given by
${\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}((\lambda^{\prime},\lambda^{\prime\prime});\mathfrak{z})$ (3.4)
@@
$:=\mathrm{Sym}_{\{x_{a}\}_{a\in(\lambda^{\prime},\lambda^{\prime\prime})}}\left(\begin{matrix}&\cr&\cr\end{matrix}\hskip-5.69054pt\prod_{a\in\lambda^{\prime},b\in\lambda^{\prime\prime}}\frac{\theta(t_{1}x^{\prime}_{a}/x^{\prime\prime}_{b})\theta(t_{2}x^{\prime}_{a}/x^{\prime\prime}_{b})}{\theta(x^{\prime}_{a}/x^{\prime\prime}_{b})\theta(\hbar x^{\prime}_{a}/x^{\prime\prime}_{b})}\right$.{}
@@
$\hskip 85.35826pt\times\begin{matrix}&\cr&\cr\end{matrix}\prod_{b\in\lambda^{\prime\prime}}{\theta(\hbar u_{1}/x_{b}^{\prime\prime})}\prod_{a\in\lambda^{\prime}}\theta(x^{\prime}_{a}/u_{2}){}$
$\left.\ \hskip 56.9055pt\begin{matrix}&\cr&\cr&\cr\end{matrix}\times{\mathrm{Stab}}_{\mathfrak{C}^{\prime},T^{1/2}}(\lambda^{\prime};\mathfrak{z}\hbar^{-1}){\mathrm{Stab}}_{\mathfrak{C}^{\prime\prime},T^{1/2}}(\lambda^{\prime\prime};\mathfrak{z})\right).{}$
Here $x_{a}^{\prime}\ (a\in\lambda^{\prime})$ and $x_{b}^{\prime\prime}\ (b\in\lambda^{\prime\prime})$ are the Chern roots for the tautological bundle on ${\mathcal{H}}_{n^{\prime}}$ and ${\mathcal{H}}_{n^{\prime\prime}}$, respectively, and we set $\{x_{a}\}_{a\in(\lambda^{\prime},\lambda^{\prime\prime})}=\{x_{a}^{\prime}\}_{a\in\lambda^{\prime}}\cup\{x_{b}^{\prime\prime}\}_{b\in\lambda^{\prime\prime}}$. Then one finds that
@@
${\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}((\lambda^{\prime},\lambda^{\prime\prime});z)$ is the elliptic stable envelope for $\mathrm{E}_{T}({\mathcal{M}}(n^{\prime}+n^{\prime\prime},2))$ with the chamber $\mathfrak{C}$ given by $|u_{1}|>|u_{2}|$. The formula (3.4) is called the shuffle product [2]. By abuse of notation we use $T^{1/2}$ in ${\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}((\lambda^{\prime},\lambda^{\prime\prime});z)$ as $T^{1/2}{\mathcal{M}}$ i.e. the polarization satisfying
$T{\mathcal{M}}=T^{1/2}{\mathcal{M}}+\hbar(T^{1/2}{\mathcal{M}})^{\vee}\in\mathrm{K}_{T}({\mathcal{M}})$
@@
for ${\mathcal{M}}={\mathcal{M}}(n^{\prime}+n^{\prime\prime},2)$. The shuffle product keeps such assignment of polarization consistent through a construction of the elliptic stable envelopes. See the remark below. In the following, we use the same convention.
@@
By repeating the shuffle product, one obtains the elliptic stable envelopes for $\mathrm{E}_{T}({\mathcal{M}}(n,r))$. Let $\lambda^{(i)}\ (i=1,\cdots,r)$ be partitions with $|\lambda^{(i)}|=n_{i}$ and consider a $r$-tuple partition ${\boldsymbol{\lambda}}=(\lambda^{(1)},\cdots,\lambda^{(r)})$. We set $|{\boldsymbol{\lambda}}|=\sum_{i=1}^{r}n_{i}=n$.
@@
The following gives the elliptic stable envelopes for $\mathrm{E}_{T}({\mathcal{M}}(n,r))$ with the chamber $\mathfrak{C}$ being $|u_{1}|>\cdots>|u_{r}|$.
@@
$\hskip-28.45274ptS^{Ell}_{\boldsymbol{\lambda}}({\bf x},{\bf u}):=\frac{\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}+1<\rho_{b}}\theta(t_{1}x_{a}/x_{b})\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}+1>\rho_{b}}\theta(t_{2}x_{b}/x_{a})}{\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}<\rho_{b}}\theta(x_{a}/x_{b})\theta(\hbar x_{a}/x_{b})}{}$
$\qquad\qquad\times\prod_{i=1}^{r}\left[\prod_{a\in{\boldsymbol{\lambda}}\atop\rho_{a}>\rho_{r_{i}}}{\theta(\hbar u_{i}/x_{a})}\prod_{a\in{\boldsymbol{\lambda}}\atop\rho_{a}\leq\rho_{r_{i}}}\theta(x_{a}/u_{i})\right],$
$\hskip-28.45274pt\sum_{{\bf t}}W^{Ell}_{T^{1/2}}({\bf t};{\bf x},{\bf u};\mathfrak{z})$
$\hskip-28.45274pt:=\hskip-14.22636pt\sum_{\delta^{(1)}\in\Upsilon_{\lambda^{(1)}},\cdots,\delta^{(r)}\in\Upsilon_{\lambda^{(r)}}}W^{Ell}_{T^{1/2}}({\bf t}^{(1)}_{\delta^{(1)}};{\bf x}^{(1)},u_{1};\mathfrak{z}\hbar^{-r+1}){}$
$\hskip 2.84544pt\times W^{Ell}_{T^{1/2}}({\bf t}^{(2)}_{\delta^{(2)}};{\bf x}^{(2)},u_{2};\mathfrak{z}\hbar^{-r+2})\cdots\times W^{Ell}_{T^{1/2}}({\bf t}^{(r)}_{\delta^{(r)}};{\bf x}^{(r)},u_{r};\mathfrak{z})$
with $r_{i}$ being the root box of $\lambda^{(i)}$, ${\bf x}^{(i)}=(x^{(i)}_{1},\cdots,x^{(i)}_{n_{i}})$ being the Chern roots associated to the Hilbert scheme ${\mathcal{H}}_{n_{i}}$ and ${\bf t}:=({\bf t}^{(1)}_{\delta^{(1)}},\cdots,\\ {\bf t}^{(r)}_{\delta^{(r)}})$.
Remark.  A more general formula for the elliptic stable envelope of affine $A$-type quivers is obtained in [3]. For ${\mathcal{M}}(n,r)$, this result coincides with Proposition 3.2. We note that the dynamical shifts of $\mathfrak{z}$ is controlled in [3] via the positive part of polarization bundle $\det T^{1/2}_{\rho>0,\lambda}$, see formula (13) and Definition 4.2.1 in [3]. For the chamber $|u_{1}|>\cdots>|u_{r}|$, these shifts are given by subtitutions $\mathfrak{z}\to\mathfrak{z}\hbar^{-r+i}$ in the functions $W^{Ell}_{T^{1/2}}$, exactly as in the formulas of Proposition 3.2.